OpenStudy (anonymous):

Factor the polynomial completely. That is, once you are finished factoring, none of the factors you obtain should be factorable: 5x^2 + 18x − 56

OpenStudy (anonymous):

http://www.algebrahelp.com/calculators/expression/factoring/ this is a good factoring calculator that walks you through the process of your factoring. Though not always comprehensively

OpenStudy (anonymous):

oh, thank you so much, ill check it out

OpenStudy (anonymous):

y = 5x^2 + 18x - 56. Use the factoring AC method Find 2 numbers b1 and b2 that: b1*b2 = ac = -280, and (b1 + b2) = 18. They are ; b1 = -10 and b2 = 28. Replace in the expression, the term 18x by the 2 terms (-10x) and (28x), then factor .

OpenStudy (anonymous):

y = 5x^2 + 18x - 56 = 5x^2 - 10x + 28x - 56 = = 5x(x - 2) + 28(x - 2) = (x - 2)(5x + 28)